Mauro Products distributes a single product, a woven basket whose selling price is $28 per unit and whose variable expense is $24 per unit. The company’s monthly fixed expense is $7,600. Required: 1. Calculate the company’s break-even point in unit sales. 2. Calculate the company’s break-even point in dollar sales. (Do not round intermediate calculations.) 3. If the company's fixed expenses increase by $600, what would become the new break-even point in unit sales? In dollar sales? (Do not round intermediate calculations.)
Answer #1

Contribution margin=Sales-Variable cost  

=(28-24)=$4 per unit

1.Breakeven=Fixed expenses/Contribution margin

=(7600/4)=1900 units

2.Breakeven=1900*28=$53200

New fixed expenses=(7600+600)=$8200

Hence new breakeven=(8200/4)=2050 units

=(2050*28)=$57400